cylindrical sample of soil is isotropically compressed under drained condition with a vertical stress of 100 kPa and a radial stress of 100 kPa. Subsequently, the axial stress was held constant and the radial stress was increased to 300 kPa under an un-drained condition.


(a) Calculate the initial mean effective stress and deviatoric stress. Create a graph with the x-axis as p, p’ and the y-axis as q (p, q space). Plot these values in (p,q) space


(b) Calculate the increase in mean total stress and deviatoric stress.


(c) Plot the total and effective stress paths (assume the soil is a linear, isotropic, elastic material).


(d) Determine the slopes of the total and effective stress paths and the maximum excess pore water pressure for each space.


1
Expert's answer
2021-06-16T08:22:01-0400

a)Initial Condition

"s'=s \frac{\sigma_a+\sigma_r}{2}=\frac{100+100}2=100kPa"

"t=\frac{\sigma_a- \sigma_r}2=\frac{100-100}2=0"

b) Loaded Condition

"\triangle s= \frac{\triangle \sigma_a+ \triangle \sigma_r}2= \frac{0+200}2=100kPa"

"\triangle s'=0=" Slope of effective stress path (ESP) for elastic soil

"\triangle t= \frac{\triangle \sigma_a- \triangle \sigma_r}2=\frac{0-200}2 kPa=-100kPa"


c)



d) Slope =1.5